Two different mixtures of fashion are available, one with 5% alcohol and the other containing 12% alcohol. How many gallons of the 12% mixture must be added to 252 gal of the 5% mixture to produce a mixture containing 9% alcohol?

Answer by checkley79(3341)

.12X+.05*252=.09(252+X)
.12X+12.6=22.68+.09X
.12X-.09X=22.68-12.6
.03X=10.08
X=10.08/.03
X=336 GAL. OF 12% ALCOHOL WAS USED.
PROOF:
.12*336+.05*252=.09(252+336)
40.32+12.60=.09*588
52.92=52.92